# Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules

### v1

@@ Acknowledgements.
- Liu et al. [2020] S. Liu, B. Cirera, Y. Sun, I. Hamada, M. Müller, A. Hammud, M. Wolf, and T. Kumagai, Dramatic enhancement of tip-enhanced Raman scattering mediated by atomic point contact formation, Nano Lett. 20, 5879 (2020).
@@
- Litman et al. [2023] Y. Litman, F. P. Bonafé, A. Akkoush, H. Appel, and M. Rossi, First-principles simulations of tip enhanced Raman scattering reveal active role of substrate on high-resolution images, J. Phys. Chem. Lette. 14, 6850 (2023).
- Pockrand and Otto [1981] I. Pockrand and A. Otto, Surface enhanced raman scattering (SERS): Annealing the silver substrate, Solid State Commun. 38, 1159 (1981).
@@
- Huber and Herzberg [1979] K. P. Huber and G. Herzberg, Molecular spectra and molecular structure: Constants of diatomic molecules (Van Nostrand, New York, 1979) pp. 716.
@@
- Ueba [1986] H. Ueba, Vibrational lineshapes of adsorbates on solid surfaces, Prog. Surf. Sci. 22, 181 (1986).
@@
- Pettinger [2010] B. Pettinger, Single-molecule surface-and tip-enhanced Raman spectroscopy, Mol. Phys. 108, 2039 (2010).
@@
The experiments were performed in an ultrahigh vacuum (UHV) chamber equipped with an STM-based TERS machine (modified UNISOKU USM-1400) at a sample temperature of 10 K. A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cleaning cycles in the chamber. H2 (99.999%, Westfalen AG) or D2 (99.7%, Air Liquide) gas was dosed into the STM unit cooled at 10 K via a variable leak valve. The gas exposure (Langmuir; L) is estimated by , where [ Torr], [s], and denote the displayed pressure value of the gauge (corrected by N2), the exposure duration, and the correction factor for the gauge. A Bayard-Alpert ionization gauge is located outside the cryogen-cooled STM unit in the chamber and (0.388) is used for H2 (D2) [59]. We used a chemically-etched Ag tip that was sharpened by focused-ion-beam milling as described elsewhere [40]. The bias voltage was applied to the sample while the tip was grounded. The STM images were obtained in the constant current mode at = 50 pA. The cleanness of the cleaned surface and tip was confirmed by STM images [Fig. S1(a)]. We also confirmed that a TERS spectrum of the Ag–Ag junction before hydrogen dosing had no molecule-derived peaks (Fig. S2, at 0 L)
In the STM junction, -polarised visible laser (532-nm solid-state or 633-nm HeNe) was incident through a fused silica window and a Ag-coated parabolic mirror (numerical aperture of 0.6). The incident laser power is 5.9 (9.7) mW at = 532 (633) nm. The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the LT-STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror. The scattered light corrected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter specific for each excitation wavelength. For the TERS and STM-induced luminescence (STML) spectra, we used the gratings 150 (for Fig. S3), 600 (Figs. 1, 2, and S2), and 1200 line/mm (Fig. 3). A sparser grating allows the acquisition of wider-wavelength range spectra, as shown in Fig. S3 (see Sec. II-B), but Raman peaks become broad and asymmetric. Thus, for the high-frequency-resolution measurements shown in Fig. 3, the densest grating was used.
@@
DFT calculations were performed with the FHI-aims program package [41, 42]. We used intermediate defaults for basis sets, radial grids and cutoff potential parameters and light defaults for the Hartree multipole expansion and angular grids. This combination proved efficient and reliable against benchmarks to tight settings. We simulated an orthogonal Ag(111) surface cell oriented perpendicular to the axis, containing four layers. The cell parameters used were Å, Å and Å. We employed a -point grid and a dipole correction was employed to maximally decouple periodic images along the direction. We used the PBE exchange-correlation functional and included vdW interactions through the pairwise screened TS-vdWsurf method of Ref. [43] in all simulations reported in the main text. Figure S4 shows a comparison of binding energies with the many-body NL-MBD dispersion correction of Ref. [60], indicating a weaker binding energy by only 5 meV and a binding distance 0.1 Å closer to the surface. Structures were optimized by allowing only hydrogen molecules and the top layer of Ag atoms of the surface to relax. All the other atoms were kept fixed at their bulk (or tip) positions. When we performed simulations including a model tip structure, this was modeled by a Ag19 cluster at varying distances from the surface. Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions close to the tip apex [46, 36].
@@
We inspected binding sites and binding energies of a single adsorbed H2 molecule, finding all values in good agreement with previous studies [44]. The binding energy was defined as the difference between the total energy when the molecule was 10 Å away from the surface and the minimum total energy of the adsorbed molecule. The blue curve in Fig. S4 shows the binding curve of a single H2 molecule adsorbed on the fcc-hollow site of the Ag(111) surface, indicating an equilibrium H2–Ag(111) bonding length of 3.09 Å and a binding energy of 32 meV. The model we used in most simulations contained 11 molecules in the unit cell (see rationale below). For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange and green curves in Fig. S4) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
@@
Performing vibrational analysis on the optimized single molecule adsorbed on the surface showed that negative frequencies related to rotational and translational modes of the molecule were always present. This led us to consider higher coverages to model local stable clustering of the molecules on the surface. We considered a full monolayer (1 ML), where all fcc-hollow sites of the unit cell were occupied by H2. With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and path-integral molecular dynamics (PIMD) at LT ( = 20 K) with the i-PI code [56] connected to FHI-aims. PIMD simulations used 12 beads in conjunction with a normal-mode colored-noise thermostat with 8 auxiliary variables and . This number of beads is not fully converged at this temperature even with the colored-noise quantum thermostats, but the purpose of these simulations was exploratory.
@@
While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate. For the quantum case we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface. We could calculate the radius of gyration of H atoms from the PIMD simulations, belonging to the 3 molecules closest to the tip at any point in the simulation. For the closest tip distance considered ( = 7.4 Å), these were 0.33 Å along and and 0.30 Å along ( = 20 K), indicating a small anisotropy. Converting these numbers to an effective thermal de Broglie wavelength and applying a inverse-square-root scaling for the temperature and mass, we would obtain Å and Å at 10 K. We note that, depending on the density of H2 molecules on the surface, is about the length of the inter-particle spacing at 10 K and signatures of bosonic exchange could begin to be observed. These features are disregarded in the current simulations.
@@
Figure S5 shows the energy profiles of a H2 molecule with varying distance to the Ag(111) surface at different tip-surface distances. We only varied the position of the molecule closest to the tip apex, while the others were kept at equilibrium [see also Figs. 3(g) and 3(i) in the main text]. The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of a H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig. S5). The asymmetry of the double-well is strongly impacted by the presence of other molecules on the surface (molecule–molecule interactions). Figure S6 shows a comparison with the limiting case of a single molecule in the unit cell at a tip–surface distance of 9.4 Å. In that case (orange curve), if the molecule can overcome the barrier, it will have a higher probability of being found closer to the tip, which was predicted in a Au-tip–single-H2–Au-surface junction [53].
I.3 C. Exact DVR Models and Vibrational Analysis
For the potentials shown in Fig. S5, we solved the one-dimensional Schrödinger equation exactly in the sinc-function discrete variable representation (DVR) [61, 62]. We set the mass of the hydrogen atom to and that of deuterium to . In all cases, calculations were performed on a regularly spaced grid of 100 points, spanning values of from to with a potential-energy cut-off . The potential energies used in the DVR were obtained from a cubic spline interpolation of the DFT-calculated values in Sec. I-B. The resulting ZPEs and average ground-state displacements of the molecular center of mass from equilibrium,
are given in Table S1, as well as those for D2. Here, is the molecular height considering the quantum mechanical ZPE contributions at the PES (i.e., the “quantum-corrected” molecular heights), and are the underlying ground-state wavefunctions, which are plotted in Fig. S7 to further illustrate the influence of tip–surface separation on the ZPEs and the equilibrium displacements.
no tip | ||||||
In parallel we conducted harmonic vibrational analysis with the i-PI code, varying the height of the H2/D2 molecule directly under the tip apex and varying the tip–sample distance [see also Figs. 3(g) and 3(i)]. Figures S9 and S9 show the dependence of the H–H and D–D stretching vibration on the mode corresponding to the vertical hindered translation of the adsorbed molecule, i.e., H2–Ag(111) and D2–Ag(111), respectively. At each , there is a red-shift of the interatomic stretching frequency with approaching tip, which follows a mass scaling between H2 and D2. As described above, with the DVR solutions, we obtain the the quantum-corrected molecule–surface distance . Based on the -dependent vibrational frequency curves, we obtain the vibrational frequency at , as shown by the blue (green) squares for H2 (D2) in Fig. S9 (S9). The plot indicates an anomalous isotopic red-shift for H2 (factor 2, instead of ), stemming from the two following factors: (1) strong anharmonic character of the hindered translation of the molecule, which is dominated by the vdW interactions with the tip and the surface (see Fig. S5), and (2) the coupling of this mode with the vibrational stretching mode.
We stress that different heights of the tip and different concentrations of the molecules can exacerbate this effect. The former because the red-shift is more pronounced at closer tip distances than at larger tip distances. The latter because weaker binding energies to the surface at lower concentrations result in softer potentials that assume a double-well character at larger tip distances (see the orange curve in Fig. S6). This effect can result in a more pronounced negative slope of the dashed curves in Figs. S9 and S9. It is challenging to determine the local molecular concentration closer to the tip apex at any measurement and a deeper study of of the vibrational frequency variation could provide indirect hints to the observed concentration. Although we neglected the anharmonic character of the H–H stretching coordinate in this study, we expect this effect not to change the qualitative discussion and introduce only rigid red-shifts on the values reported here.
@@
We conducted TERS simulations according to the methodology proposed in Ref. [46]. We used Tip A discussed in the reference and calculated the anharmonic Raman spectrum based on a classical ab initio molecular dynamics (AIMD) trajectory (5 ps) of a molecular layer adsorbed on the surface and a tip apex located 4 Å above the molecular layer at = 20 K. We calculated the TERS intensity as , where was obtained by a local polarizability calculation at each geometry in the time-series given by the molecular dynamics. For this part of the calculation we only considered the free-standing molecular layer (no surface) at the positions adopted in the AIMD trajectory including the surface. We note that the method of Ref. [46] is unable to capture the electromagnetic plasmonic enhancement in its totality. We do not observe any noteworthy features in the spectrum and it overall agrees with the experimental spectrum. The red-shift of the rotational band in the simulation is explained by the lack of consideration of the nuclear spin, making calculations reproduce only classical rotations that are red-shifted if compared to the quantized transition of para-H2. This simulation had only an exploratory character and achieving statistical convergence through the use of parametrized machine-learned potentials, plus the inclusion of accurate plasmonic enhancements is beyond the scope of this paper.
@@
To evaluate the electric field in the STM junction, we performed finite element method (FEM) simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module). We used the three-dimentional model with a Ag tip and Ag sample plate described in Supporting Information of Ref. [51]. As needed, at the apex of the Ag tip (a length of 300 nm, an effective radius of 30 nm), a half-sphere protrusion with a radius of 0.5 nm is attached as a “nanotip.” A -polarized plane wave with of 532 nm and an electric field of 1 V/m is incident with an angle of 35∘ relative to the sample surface. The perfectly matched layers surrounding the volume are set to absorb all outgoing waves. The dielectric constant of the Ag objects is referenced from the literature [63]. The surface normal component of the electric field is sampled.
A Ag flat plate is placed beneath the tip apex with a gap distance . Here we define the origin of the gap distance ( = 0) as the tip height when the Ag-tip apex forms point contact with the Ag surface. As described above, the quantum-corrected molecular height is calculated to be approximately 0.32–0.33 nm for D2/Ag(111). In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model. Although the calculated depends on the tip height [Figs. 3(h) and 3(j) in the main text], in the FEM model, we used the constant height for the sampling point at any gap distances. We confirmed that differences in the definition or the molecular position on the order of 0.1 nm do not result in qualitative differences in the field enhancement factor curves. To directly compare with the experimental TERS intensities, the electric field enhancement factor (Ref. [64]) is displayed in the plots of Fig. 2(c) in the main text. To the field enhancement of the nearest neighboring molecule, assuming the molecule is placed in the next adsorption site of Ag(111), = 0.29 nm is used [red triangle in the inset of Fig. 2(c) in the main text]. Thus, when the point at the surface just under the tip is defined as (, , ) = (0, 0, 0), the black and red curves shown in Fig. 2c of the main text represent the plots sampling at (, , ) = (0, 0, 0.20) nm and (0.29, 0, 0.20) nm, respectively.
The tip heights in the DFT calculations are comparable with the gap distances with the FEM simulations that reproduced the Raman enhancement [Fig. 2(c) in the main text; the gap distances of 550 to 850 pm in the simulation correspond to the tip height of 0.8 to 1.1 nm, considering the Ag-atom radius].
@@
H2 and D2 have two nuclear spin isomers, ortho and para. para-H2 (nuclear spin quantum number = 0) is more stable than ortho-H2 ( = 1) by 14 meV [65]. para- (ortho-) H2 has only even (odd) rotational quantum numbers . Although ortho-H2 is the predominant species in the gas phase at room temperature (a temperature equilibrium composition of 75%), the molecular adsorption on metals enhances the ortho-to-para conversion, leading to the dominance of the para species on the surfaces [4]. In the TERS spectra, we can detect only -H2 (Table S2); as a reference, in the gas phase, ortho-H2 have a Raman-active rotational mode [] at 587 cm-1 [65, 48]. This implies that we detected only the physisorbed species after ortho-para conversion on the surface. In the case of D2, ortho ( = 0, 2) is more stable than para ( = 1) by 7.4 meV and more dominant (67%) in the gas phase at room temperature. Only -D2 was detected in our TERS spectra, indicating again that para-D2 was converted on the surface.
@@
Although the dominance of para-H2 in the TERS spectra is consistent with a previous HREELS study of H2 on a Ag film (0.5–10 L) [8], some previous reports detect ortho-H2 on Ag(111). A HREELS study on H2/Ag(111) [66] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr). Photostimulated desorption studies (30 L) [12, 13] shows that ortho-H2 and para-D2 are alive on Ag(111) on the order of hundred seconds and that the conversion of D2 is slower than H2.
@@
To compare with the above-mentioned studies, we performed TERS measurements during exposing D2 to a clean Ag(111) surfaces, as shown in Fig. S2. The gas pressure was set to Torr, comparable with the EELS study [66]. The ortho-D2-derived peaks appeared at 16 L and then saturated at 40 L, while para-D2 was not detected. The weak intensity at low exposures may be due to molecular diffusion along the surface at low coverages and the weak contribution of neighboring molecules to the TERS signal [see Fig. 2(c) in the main text]. It should be noted that even if the gas pressure and exposure values are comparable, the molecular coverage varies depending on the chamber configuration. Nevertheless, we confirmed that ortho-H2 and para-D2 were not detected at any pressures and exposures we used (0.8–2 Torr; 102–103 Langmuir). The HREELS observation of the metastable isomer only at high exposures [66] may be attributed to be hydrogen multilayers, where the second layer does not interact with the Ag surface. In contrast, the narrow TERS gap for the LT-TERS measurement would suppress the formation of the multiple layers. Another possibility that the metastable isomer were not detected in TERS is that the LSPR in the gap is accelerating the para-to-ortho conversion of the molecule under the tip. Although far-field 532-nm irradiation did not affect the conversion rate [13], the strong electric field localized in the STM junction could interact with the nuclear spin of the adsorbate. This effect would be verified through theoretical calculations.

### v2

@@ Acknowledgements.
- Sugimoto and Fukutani [2014] T. Sugimoto and K. Fukutani, Effects of rotational-symmetry breaking on physisorption of ortho-and para-H2 on Ag(111), Phys. Rev. Lett. 112, 146101 (2014).
@@
- Merino et al. [2018] P. Merino, A. Rosławska, C. C. Leon, A. Grewal, C. Große, C. González, K. Kuhnke, and K. Kern, A single hydrogen molecule as an intensity chopper in an electrically driven plasmonic nanocavity, Nano Lett. 19, 235 (2018).
@@
- Haran [2010] G. Haran, Single-molecule Raman spectroscopy: a probe of surface dynamics and plasmonic fields, Acc. Chem. Res. 43, 1135 (2010).
@@
- Park et al. [2024] Y. Park, I. Hamada, A. Hammud, T. Kumagai, M. Wolf, and A. Shiotari, Atomic-precision control of plasmon-induced single-molecule switching in a metal–semiconductor nanojunction, Nat. Commun. 15, 6709 (2024).
@@
- Litman et al. [2023] Y. Litman, F. P. Bonafé, A. Akkoush, H. Appel, and M. Rossi, First-principles simulations of tip enhanced Raman scattering reveal active role of substrate on high-resolution images, J. Phys. Chem. Lett. 14, 6850 (2023).
@@
- Cirera et al. [2022] B. Cirera, Y. Litman, C. Lin, A. Akkoush, A. Hammud, M. Wolf, M. Rossi, and T. Kumagai, Charge transfer-mediated dramatic enhancement of Raman scattering upon molecular point contact formation, Nano Lett. 22, 2170 (2022).
- Cui et al. [2023] J. Cui, Y. Zhang, and Z.-C. Dong, Atomistic polarization model for Raman scattering simulations of large metal tips with atomic-scale protrusions at the tip apex, Nanophotonics 12, 4031 (2023).
@@
- Pettinger et al. [2009] B. Pettinger, K. F. Domke, D. Zhang, G. Picardi, and R. Schuster, Tip-enhanced Raman scattering: influence of the tip-surface geometry on optical resonance and enhancement, Surf. Sci. 603, 1335 (2009).
@@
- Markland et al. [2008] T. E. Markland, S. Habershon, and D. E. Manolopoulos, Quantum diffusion of hydrogen and muonium atoms in liquid water and hexagonal ice, J. Chem. Phys. 128, 194506 (2008).
- Megaw and Simon [1936] H. D. Megaw and F. Simon, Density and compressibility of solid hydrogen and deuterium at 4.2 K, Nature 138, 244 (1936).
@@
The experiments were performed in an ultrahigh vacuum (UHV) chamber equipped with an STM-based TERS machine (modified UNISOKU USM-1400) at a sample temperature of 10 K. A single-crystalline Ag(111) surface (MaTeck) was cleaned by Ar-sputtering-and-annealing cycles in the chamber. H2 (99.999%, Westfalen AG) or D2 (99.7%, Air Liquide) gas was introduced into the STM unit cooled at 10 K via a variable leak valve. The gas exposure (Langmuir; L) is estimated by , where [ Torr], [s], and denote the displayed pressure value of the gauge (corrected by N2), the exposure duration, and the correction factor for the gauge, respectively. A Bayard-Alpert ionization gauge is located outside the cryogen-cooled STM unit in the chamber and we used for H2 and for D2 [63]. We used a chemically-etched Ag tip that was further sharpened by focused-ion-beam milling as described elsewhere [40]. The bias voltage was applied to the sample while the tip was grounded. The STM images were obtained in the constant current mode at = 50 pA. The cleanness of the Ag(111) surface and tip was confirmed by STM images [Fig. S1(a)]. We also confirmed that a TERS spectrum of the Ag–Ag junction before hydrogen dosing had no molecule-derived peaks (Fig. S2, at 0 L).
In the STM junction, -polarised visible laser (532-nm solid-state or 633-nm HeNe) was directed through a fused silica window and an Ag-coated parabolic mirror (numerical aperture of 0.6). The incident laser power is 5.9 mW at = 532 nm and 9.7 mW at 633 nm. The parabolic mirror is mounted on a five-axis piezoelectric-motor stage in the LT-STM unit and the laser is focused at the tip apex by adjusting the position and angle of the parabolic mirror. The scattered light collected by the parabolic mirror was directed to a spectrometer (AndorShamrock 303i) outside the UHV chamber via a beam splitter and a longpass filter, each selected according to the excitation wavelength. For the TERS and STM-induced luminescence (STML) spectra, we used the gratings 150 (for Fig. S3), 600 (Figs. 1, 2, and S2), and 1200 line/mm (Fig. 3 and S10). A lower-density grating allows the acquisition of spectra over a wider wavelength range, as shown in Fig. S3 (see Sect. II-B), but resulting in broad and asymmetric Raman peaks. Therefore, for the high-frequency-resolution measurements shown in Fig. 3, the densest grating was used.
@@
DFT calculations were performed with the FHI-aims program package [49, 50]. We used intermediate defaults for basis sets, radial grids and cutoff potential parameters and light defaults for the Hartree multipole expansion and angular grids. This combination proved efficient and reliable against benchmarks to tight settings. We simulated an orthogonal Ag(111) surface cell oriented perpendicular to the axis, containing four layers. The cell parameters used were Å, Å and Å. We employed a -point grid and a dipole correction was employed to maximally decouple periodic images along the direction. We used the PBE exchange-correlation functional and included vdW interactions through the pairwise screened TS-vdWsurf method of Ref. [51] in all simulations reported in the main text. Figure S4 shows a comparison of binding energies with the many-body MBD-NL dispersion correction of Ref. [64], indicating a weaker binding energy by only 5 meV and a binding distance 0.1 Å closer to the surface. Structures were optimized by allowing only hydrogen molecules and the top layer of Ag atoms of the surface to relax. All the other atoms were kept fixed at their bulk (or tip) positions. When we performed simulations including a model tip structure, this was modeled by an Ag19 cluster at varying distances from the surface. The geometry of the tip is shown in Fig. S5. Although this tip size would not be sufficient to capture a proper plasmonic response of the tip, the tip structure can capture the local interactions with molecules close to the tip apex [46, 54, 41].
@@
We inspected binding sites and binding energies of a single adsorbed H2 molecule, finding all values in good agreement with previous studies [52]. The binding energy was defined as the difference between the total energy when the molecule was 10 Å away from the surface and the minimum total energy of the adsorbed molecule. The blue curve in Fig. S4 shows the binding curve of a single H2 molecule adsorbed on the fcc-hollow site of the Ag(111) surface, indicating an equilibrium H2–Ag(111) bonding length of 3.09 Å and a binding energy of 32 meV. The model we used in most simulations contained 11 molecules in the unit cell (see rationale below). For this molecular concentration, we evaluated the intermolecular interaction by comparing the binding curves (orange curve in Fig. S4) with that of the single-molecule case (blue); the molecule–molecule interaction in our models amounts to 55 meV/molecule.
@@
Performing vibrational analysis on the optimized single molecule adsorbed on the surface showed that negative frequencies related to rotational and translational modes of the molecule were always present. This led us to consider higher coverages to model local stable clustering of the molecules on the surface. We considered a full monolayer (1 ML), where all fcc-hollow sites of the unit cell were occupied by H2. With this setup we performed Born-Oppenheimer molecular dynamics (BOMD) and PIMD at LT ( = 20 K) with the i-PI code [65] connected to FHI-aims. PIMD simulations used 12 beads in conjunction with a normal-mode colored-noise thermostat with 8 auxiliary variables and . This number of beads is not fully converged at this temperature even with the colored-noise quantum thermostats, but the purpose of these simulations was exploratory.
@@
While the full monolayer was stable during more than 10 ps on the classical trajectories, in the quantum trajectory 5 molecules readily desorbed (within the first picosecond) from the substrate. For the quantum case, we validated this observation by simulating trajectories with a parameterized Lennard-Jones potential representing the Ag(111) surface. We could calculate the radius of gyration of H atoms from the PIMD simulations, belonging to the 3 molecules closest to the tip at any point in the simulation. For the closest tip distance considered ( = 7.4 Å), these were 0.33 Å along and and 0.30 Å along ( = 20 K), indicating a small anisotropy. Converting these numbers to an effective thermal de Broglie wavelength and applying a inverse-square-root scaling for the temperature and mass, we would obtain Å for H2 and Å for D2 at 10 K. We note that, depending on the density of H2 molecules on the surface, for H2 is about the length of the inter-particle spacing at 10 K and signatures of bosonic exchange could begin to be observed. These features are disregarded in the current simulations. Further, we note that even though these simulations provide a clear indication that quantum mechanical effects affect the density of the molecular coverage of the surface, a complete determination would require much longer simulations, which would take years to complete or require the development of a tailored machine-learning potential that surpasses the scope of this work.
@@
Figure S6 shows the energy profiles experienced by an H2 molecule when varying the distance of its center of mass with respect to the Ag(111) surface, at different tip-surface distances, and at 0.69 ML coverage. We only varied the position of the molecule closest to the tip apex, while the others were kept at equilibrium. The profile shapes are similar to an asymmetric double-well potential at closer tip–surface distances, while they assume a shape closer to a Morse potential at larger distances, which asymptotically approaches the profile of an H2 molecule on Ag(111) without the Ag tip (yellow curve in Fig. S6). The asymmetry of the double-well is strongly impacted by the presence of other molecules on the surface (molecule–molecule interactions). Figure S7(a) shows a comparison with the limiting case of a single molecule in the unit cell (0.06 ML coverage) at a tip–surface distance of 9.4 Å. In that case (orange curve), if the molecule can overcome the barrier, it will have a higher probability of being found closer to the tip, which was predicted in a Au-tip–single-H2–Au-surface junction [17]. As we discuss in the main text, this effect is essential to explain the isotope-dependent red-shift observed experimentally.
Figure S7(b) shows a comparison of the 0.69 ML coverage with the tip fixed at Å and different vdW dispersion corrections, added to the PBE functional. It is clear that details of the vdW dispersion impact the shape of the potential. In this case, the many-body vdW dispersion decreases the distance between the two minima of the asymmetric double-well
I.3 C. Model 2D Schrödinger Equation
[…]
where is the potential energy of a molecule with its center of mass fixed at a distance above the substrate surface and with all other coordinates kept at their equilibrium values, as described at the end of Sec. I.2. The vibrational motion is parametrized by the reduced mass and angular frequency . We set the mass of a hydrogen atom to and use .
The vibrational frequencies are computed from numerical (finite-difference) second derivatives of the DFT energies with respect to the vibrational coordinate,
[…]
We computed the numerically exact solutions to the 2D Schrödinger equation corresponding to Eq. (S1) using the sinc-function discrete variable representation (DVR) [66, 67]. The equation was solved for high coverage (0.69 ML) with and . Given the large separation in the timescales of translational and vibrational motion, we also considered the vibrationally adiabatic approximation, whereby one fixes the translational coordinate in Eq. (S1) and solves the corresponding one-dimensional (1D) Schrödinger equation in , yielding the analytical harmonic vibrational energies
These are then used to construct the vibrationally adiabatic translational potential,
for which we solve the 1D Schrödinger equation numerically using DVR. The energy difference between the ground state solutions for and is an excellent approximation to the vibrational excitation energy of the 2D model (see Table S1). Given the accuracy of the vibrationally adiabatic approximation, we adopt it in all subsequent calculations.
2D | ||||
1D |
[…]
11.4 | ||||||
9.4 | ||||||
7.9 | ||||||
7.4 |
11.4 | ||||||
9.4 | ||||||
7.9 | ||||||
7.4 |
@@
We conducted TERS simulations according to the methodology proposed in Refs. [46]. We used Tip A discussed in Ref. [54] and calculated the anharmonic Raman spectrum based on an ab initio molecular dynamics (AIMD) trajectory (5 ps) of a molecular layer adsorbed on the surface and a tip apex located 4 Å above the molecular layer at = 20 K. We calculated the TERS intensity as , where was obtained by a local polarizability calculation at each geometry in the time-series given by the molecular dynamics. For this part of the calculation, we only considered the free-standing molecular layer (no surface) at the positions adopted in the AIMD trajectory including the surface. We note that the method of Ref. [46] is unable to capture the plasmonic enhancement in its totality. We do not observe any noteworthy features in the calculated spectrum and it overall agrees with the experimental spectrum. The red-shift of the rotational band in the simulation is explained by the absence of nuclear spin in the simulations, making calculations reproduce only classical rotations that are red-shifted compared to the quantized transition of para-H2. This simulation is of exploratory character, as achieving statistical convergence through the use of parametrized machine-learned potentials along with accurate plasmonic enhancements is beyond the scope of this paper.
@@
To evaluate the electric field in the STM junction, we performed FEM simulations using COMSOL Multiphysics (version 6.2 with the Wave Optics Module). We used the three-dimentional model with an Ag tip and Ag sample plate described in Supporting Information of Ref. [40]. To reproduce a plasmonic picocavity in the STM junction, at the apex of the Ag tip (a length of 300 nm, an effective radius of 30 nm), an Ag half sphere with a radius of 0.5 nm is attached as an atomic-scale protrusion. For comparison, we also constructed another Ag tip model with an effective radius of 30 nm without the atomic-scale protrusion [see the orange dotted curve in Fig. 2(c) in the main text]. A -polarized plane wave with of 532 nm and an electric field of 1 V/m is incident with an angle of 55∘ relative to the sample normal. The perfectly matched layers surrounding the volume are set to absorb all outgoing waves. The dielectric constant of the Ag objects is referenced from the literature [68]. The surface normal component of the electric field is sampled.
An Ag flat plate is placed beneath the tip apex with a gap distance . Here we define the origin of the gap distance ( = 0) as the tip height when the Ag-tip apex forms point contact with the Ag surface. As shown in Table S3, the molecular height is calculated to be 0.32–0.33 nm for D2/Ag(111). In the definition, taking into account the atomic radius of Ag, the position of the molecule corresponds to 0.2 nm above the Ag surface plane in the FEM model. Although the calculated depends on the tip height (Table S3), in the FEM model, we used the constant height for the sampling point at any gap distances. We confirmed that differences in the definition or the molecular position on the order of 0.1 nm do not result in qualitative differences in the field enhancement factor curves. To directly compare with the experimental TERS intensities, the electric field enhancement factor (Ref. [57]) is displayed in the plots of Fig. 2(c) in the main text. To the field enhancement of the nearest neighboring molecule, assuming the molecule is placed in the next adsorption site of Ag(111), = 0.29 nm is used [red triangle in the inset of Fig. 2(c) in the main text]. Thus, when the point at the surface just under the tip is defined as (, , ) = (0, 0, 0), the black and red curves shown in Fig. 2c of the main text represent the plots sampling at (, , ) = (0, 0, 0.20) nm and (0.29, 0, 0.20) nm, respectively.
As shown in Fig. 2, the gap distances with the FEM simulations pm to 850 pm well reproduced the Raman enhancement at the experimental tip height to 0 pm. Considering the Ag-atom radius, for the simulations (Fig. 4 in the main text) is comparable to nm. Therefore, pm corresponds to nm, which was used for the alignment of the experimental and theoretical plots displayed in Fig. 4(e) of the main text.
@@
H2 and D2 have two nuclear spin isomers, ortho and para. para-H2 (nuclear spin quantum number = 0) is more stable than ortho-H2 ( = 1) by 14 meV [69]. para-H2 has only even rotational quantum numbers , whereas and ortho-H2 has only odd . Although ortho-H2 is the predominant species in the gas phase at room temperature (a temperature equilibrium composition of 75%), the molecular adsorption on metals enhances the ortho-to-para conversion, leading to the dominance of the para species on the surfaces [4]. In the TERS spectra, we can detect only -H2 (Table S4); as a reference, in the gas phase, ortho-H2 have a Raman-active rotational mode [] at 587 cm-1 [69, 56]. This implies that we detected only the physisorbed species after ortho-para conversion on the surface. In the case of D2, ortho ( = 0, 2) is more stable than para ( = 1) by 7.4 meV and more dominant (67%) in the gas phase at room temperature. Only -D2 was detected in our TERS spectra, indicating again that para-D2 was converted on the surface.
@@
Although we introduced normal hydrogen gases (n-H2 and n-D2) into the chamber, we observed rotational/vibrational transitions only for para-H2 and ortho-D2 in the TERS spectra (Fig. 1 in the main text). The dominance of para-H2 on Ag(111) in the TERS spectra ( = 10 K) is consistent with a previous HREELS study of H2 on an Ag film (exposures of 0.5–10 L, = 10 K) [8], whereas some previous reports detect ortho-H2 on Ag(111). A HREELS study on H2/Ag(111) at 6 K [11] shows that rotational mode peaks of both para-H2 and ortho-H2 were detected under a higher H2 pressure ( Torr) while para-H2 was dominantly detected under a lower pressure ( Torr). Photostimulated desorption studies [14, 15] shows that at an exposure of 30 L and a temperature of 6 K, ortho-H2 and para-D2 are alive on Ag(111) on the order of hundred seconds and that the conversion of D2 is slower than H2.
@@
To compare with the above-mentioned studies, we performed TERS measurements while exposing D2 to a clean Ag(111) surfaces, as shown in Fig. S2. The gas pressure was set to Torr, comparable with the EELS study [11]. The ortho-D2-derived peaks appeared at 16 L and then saturated at 40 L, while para-D2 was not detected. The weak intensity at low exposures may be due to molecular diffusion along the surface at low coverages and the weak contribution of neighboring molecules to the TERS signal [see Fig. 2(c) in the main text]. It should be noted that even if the gas pressure and exposure values are comparable, the molecular coverage varies depending on the chamber configuration. Nevertheless, we confirmed that ortho-H2 and para-D2 were not detected at any pressures and exposures we used (0.8–2 Torr; 102–103 L). The HREELS observation of the metastable isomer at 6 K only at high exposures [11] may be attributed to be hydrogen multilayers, where the second layer does not interact with the Ag surface. For our LT-TERS measurement, in contrast, only a monolayer would be formed at the higher temperature, 10 K. Another possibility that the metastable isomer were not detected in TERS is that the LSPR in the gap is accelerating the para-to-ortho conversion of the molecule under the tip. Although far-field 532-nm irradiation did not affect the conversion rate [15], the strong electric field localized in the STM junction could interact with the nuclear spin of the adsorbate.